OpenStudy (anonymous):

Solve the equation: x2 - 12x = -36

OpenStudy (anonymous):

\[(x - 6)^2 = 0\] \[x = 6\]

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

(a-b)^2=a^2-2ab+b^2 in our case: a is x, b is 6

OpenStudy (anonymous):

x2-12x=-36 x2-12x+36=0 x2-6x-6x+36=0 x(x-6)-6(x-6)=0 (x-6)(x-6) x=0